What length of copper wire 1.6 mm in diameter is needed to make a 18 ohm resistor at 18 degrees C.
Answer #1

i am taking resistivity  of copper in this question ρ = 1.724 x 10-8

if you have different data , then let me know

  

R = ρ L /A

A = pi*r^2 = pi* ( 1.6*10^-3 /2)^2 =2.0106193e-6 m2

R = ρ L /A

18 = 1.724 x 10-8 *L /(2.0106193e-6)
from here L≈2099.25 m

so length we need L ≈2099.25 m

***************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..